OpenStudy (anonymous):

which point lies in the solution set for the system: 2y-x>or equal to -6 2y-3x<-6

OpenStudy (anonymous):

Graphs the lines: 2y-x=-6 and 2y-3x =-6 They intersects at (0.-3). 1st inequality says it is "above" the 2y-x=-6 line 2nd says it is "below" the 2y-3x=-6 line Shade in the area between these lines => that shaded part is the solution set

OpenStudy (anonymous):

thank you again!!!!